OpenStudy (anonymous):

Historical Feature: In the equation x^3+px+q=o, replace x by H+K. Let 3HK=-p, and show that H^3+K^3=-q

OpenStudy (perl):

this is how you solve the cubic equation

OpenStudy (perl):

In the equation $$ \Large{ x^3+px+q=0 \\ x= h+k \\ 3hk=-p \\~\\ \textrm{Goal: Show that} ~h^3+k^3=-q } $$